# Boundary Attention: Learning to Find Faint Boundaries at Any Resolution

### 01 Jan 2024 v1

@@ 6 Conclusion
We have introduced a differentiable model that explicitly reasons about geometric primitives such as edges, corners, junctions, and regions of uniform appearance, by using boundary attention. Despite being trained on simple geometric synthetic data, our method generalizes to natural images, and predicts clean boundaries, even when the images it is applied to are significantly corrupted by noise. Furthermore, the parametric form of our model elicits a natural formulation for boundaries with subpixel precision, and its bottom-up patch-wise approach enables it to process any image at its native resolution and aspect ratio.
Supplemental Material
Boundary Attention: Learning to Find Faint Boundaries at Any Resolution
1 . The space of $M$-junctions
@@
Figure 11: Anatomy of an $M$-junction $\mathbf{g}=(\mathbf{u},\theta,\boldsymbol{\omega})$ with $M=3$. Left: Boundary directions $\phi_{j}$ and central directions $\psi_{j}$ are determined directly from relative angles $\boldsymbol{\omega}$ and orientation $\theta$ (which is equal to $\phi_{1}$). Middle panels: Unsigned distance function for a boundary ray $d_{3}(x;\mathbf{g})$ and overall unsigned distance function $d(x;\mathbf{g})$, which is the minimum of the three per-ray ones. Right: Associated boundary function $b_{\eta}(x;\mathbf{g})$ using $\eta=0.7$.
As shown in Figure 11, other geometric features of a junction can be directly derived from the orientation and angles. The central directions $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{M})$ are
@@
Figure 11 shows an example of a junction’s distance function and its associated boundary function with $\eta=0.7$.
Interpolation
@@
2 . Training Data
Figure 12: Columns 1 to 5: Examples of the synthetic data used to train our model using supervision with ground-truth boundaries. Column 6: Rendered distance maps corresponding to column 5. The training data contains random circles and triangles that each have a random RGB color, and the images are corrupted by various types and amounts of noise. Each noiseless image has an unrasterized, vector-graphics representation of its shapes and colors, which specify the clean image and exact boundary-distance map with unlimited resolution.
We find that we can train our model to a useful state using purely synthetic data, examples of which are depicted in Figure 12. In fact, we find it sufficient to use very simple synthetic data that consists of only two basic shapes—circles and triangles—because these can already produce a diverse set of local edges, thin bars, curves, corners, and junctions, in addition to uniform regions. We generate an image by randomly sampling a set of circles and triangles with geometric parameters expressed in continuous, normalized image coordinates $[0,1]\times[0,1]$. We then choose a random depth ordering of the shapes, and we choose a random RGB color for each shape. Importantly, the shape and color elements are specified using a vector-graphics representation, and the shape elements are simple enough to provide an exact, symbolic expression for each image’s true boundary-distance map, without approximation or rasterization. They also allow calculating the precise locations, up to machine precision, for all of the visible corners and junctions in each image.
@@
S3.2 . Neighborhood Cross-attention
@@
The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block. We use $100,000$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block. Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes).
Our trained weights will be released upon publication.
4 . Qualitative Behavior for Natural Images
In Figures 14 and 13, we show how the model behaves on noiseless natural images that contain texture and recognizable objects. In particular, Figure 13 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations. The figure compares our output to that from Canny [4], Field of Junctions [33] with a patch size of $11$, Pidinet [31], and EDTER [28]. The latter two methods are trained on human annotated data, whereas the former two methods, like our model, are not. (Note that inputs for all models besides EDTER [28] were $300\times 400$. Input to EDTER was down-sampled to $225\times 300$ due to its input size constraint.)
We find that our model produces finer structures than the end-to-end learned models [10, 28] because it is trained to only use local spatial averages of color as its cue for boundaries and grouping. It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans. Compared to the bottom-up methods of Canny [4] and Field of Junctions [33], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of field $\mathbf{p}[n]$. In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
@@
Figure 13: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [28] and Pidinet [31] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [4], and Field of Junctions [33] with patch size $11$.
@@
Figure 14: Qualitative behavior of our model on noiseless natural images. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
5 . Additional Examples for Low-light Images
@@
Figure 15 shows examples of applying our model to indoor images taken by an iPhone XS in low light conditions.
Figure 15: Visualization of our model’s output for low-light images captured by an iPhone XS. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
@@
Figure 16 provides additional comparisons for a sample of varying-noise images from the ELD dataset [34]. When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases. Some algorithms, such as Field of Junctions [33], have tunable parameters such as patch-size that provide control over the level of detection. A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels. Conversely, a large patchsize provides more resilience to noise but has not ability to recover fine structure at all. Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
@@
In Figure 16 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods. We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve. We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions. Increasing the local size improves their resilience to noise but reduces their spatial precision. Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
Figure 16: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise. We compare to end-to-end models that are trained to match human annotations (EDTER [28], HED [35], Pidinet [31], and Structured Forests [10]) in addition to low-level models that are not (Canny [4], and Field of Junctions (FOJ) [33]).
@@
Figure 20 contains additional examples of images cropped from the ELD dataset. Here we include examples with even higher levels of noise to show the complete degradation of our algorithm and others.
Figure 17:
Figure 18: (cont.)
Figure 19: (cont.)
@@
Figure 20: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
6 . Additional Uses of Our Model
@@
S6.1 . Color-based Depth Completion
Figure 21 shows an example of using our model for simple hole-filling in the depth channels of RGBD images from the Middlebury Stereo Datasets [30, 29]. We run our model on the RGB channels, and then for each pixel $n$ that has a missing depth value, we use our model’s output local attention kernels $a_{n}(x)$ to fill in that pixel’s value using an attention-weighted average of the observed depth values around it. This simple algorithm can be applied whenever the hole sizes are smaller than the maximum diameter of our attention maps, which is $34\times 34$ in our current implementation).
Figure 21: Using our model for depth completion in RGBD images. Left: Input RGB channels. Middle: Input depth channel, with dark blue indicating missing values. Right: Completed depth using our model’s output attention kernels.
S6.2 . Application: Photo Stylization
Figure 22 shows examples of using our model’s output for image stylization, by superimposing an inverted copy of the output boundary map $\bar{b}_{\eta}[n]$ onto the smoothed colors $\bar{\mathbf{f}}[n]$.
Figure 22: Examples of stylized natural photographs, created by imposing our method’s output boundary map onto the output smoothed colors.

### 18 Mar 2024 v2

@@ 7 Conclusion
We have introduced a differentiable model that uses boundary attention to explicitly reason about geometric primitives such as edges, corners, junctions, and regions of uniform appearance in images. Our bottom-up, feedforward network can encode images of any resolution and aspect ratio into a field of geometric primitives that describe the local structure within every patch. Despite being trained on simple geometric synthetic data, our method can consistently localize boundaries under high levels of noise better than state-of-the-art deep networks, while achieving similar performance to recent optimization-based approaches in a fraction of the time.
Supplemental Material for Boundary Attention: Learning to Localize Boundaries under High Noise
S1 The space of $M$-junctions
@@
Fig. S1: Anatomy of an $M$-junction $\mathbf{g}=(\mathbf{u},\theta,\boldsymbol{\omega})$ with $M=3$. Left: Boundary directions $\phi_{j}$ and central directions $\psi_{j}$ are determined directly from relative angles $\boldsymbol{\omega}$ and orientation $\theta$ (which is equal to $\phi_{1}$). Middle panels: Unsigned distance function for a boundary ray $d_{3}(x;\mathbf{g})$ and overall unsigned distance function $d(x;\mathbf{g})$, which is the minimum of the three per-ray ones. Right: Associated boundary function $b_{\eta}(x;\mathbf{g})$ using $\eta=0.7$.
As shown in Figure S1, other geometric features of a junction can be directly derived from the orientation and angles. The central directions $\boldsymbol{\psi}=(\psi_{1},\ldots,\psi_{M})$ are
@@
Figure S1 shows an example of a junction’s distance function and its associated boundary function with $\eta=0.7$.
S1.1  Interpolation
@@
S2 Training Data
Fig. S2: Columns 1 to 5: Examples of the synthetic data used to train our model using supervision with ground-truth boundaries. Column 6: Rendered distance maps corresponding to column 5. The training data contains random circles and triangles that each have a random RGB color, and the images are corrupted by various types and amounts of noise. Each noiseless image has an unrasterized, vector-graphics representation of its shapes and colors, which specify the clean image and exact boundary-distance map with unlimited resolution.
We find that we can train our model to a useful state using purely synthetic data, examples of which are depicted in Figure S2. In fact, we find it sufficient to use very simple synthetic data that consists of only two basic shapes—circles and triangles—because these can already produce a diverse set of local edges, thin bars, curves, corners, and junctions, in addition to uniform regions. We generate an image by randomly sampling a set of circles and triangles with geometric parameters expressed in continuous, normalized image coordinates $[0,1]\times[0,1]$. We then choose a random depth ordering of the shapes, and we choose a random RGB color for each shape. Importantly, the shape and color elements are specified using a vector-graphics representation, and the shape elements are simple enough to provide an exact, symbolic expression for each image’s true boundary-distance map, without approximation or rasterization. They also allow calculating the precise locations, up to machine precision, for all of the visible corners and junctions in each image.
@@
S3.2  Neighborhood Cross-attention
@@
The final stage of training adds a second boundary attention block with weights that are initialized using a copy of the pretrained weights of the first boundary attention block. We use $100k$ crops of size $125\times 125$ from our Kaleidoshape images (10% withheld for testing) and the full set of losses; and we optimize all of the model’s parameters, including those of the neighborhood MLP-mixer and the first boundary attention block. Like in pretraining, we add a spatial importance that prioritizes region containing a corner (i.e., a visible triangle vertex) or a junction (i.e., a visible intersection between the boundaries of any two shapes). Our trained weights will be released upon publication.
S4 Model Behavior
S4.1  Generalization to Unseen Shapes
During training, our model observes only circles and triangles of varying sizes and colors. We measure how well our model generalizes to unseen shapes by creating a new synthetic dataset containing rectangles, polygons, and ellipses of various sizes, colors, and rotations. We evaluate our model against low-level methods as well as methods trained on semantic data, and note the results here closely align with the performance trends on kaleidoshapes (Fig. 5), suggesting that our bottom-up model can generalize to unseen geometries.
Fig. S3: We measure how well our model can generalize to shapes outside of the training distribution by evaluating our model and others on noisy synthetic data composed exclusively of unseen shapes.
S4.2  Linear Interpolation in Junction Space
Surprisingly, we note that our network learns a spatially smooth manifold of junctions in its hidden state. Figure S4 visualizes some properties of its learned embedding of junction space. We generate equally-spaced samples $\boldsymbol{\gamma}_{i}\in\mathbb{R}^{D_{\boldsymbol{\gamma}}}$ by linearly interpolating from a particular $\boldsymbol{\gamma}_{a}$ to 00 and then to a particular $\boldsymbol{\gamma}_{b}$. We see that the embedding space is smooth, and interestingly, that it learns to associate its zero with nearly-equal angles and a vertex close to the patch center. For visual comparison, we show an analogous interpolation from $\mathbf{g}_{a}$ to $\mathbf{g}_{0}\triangleq(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$ and then to $\mathbf{g}_{b}$ by using a linear geometric interpolation in junction space ${\cal G}$, the expressions for which are in the supplement.
Fig. S4: Top: Linear interpolation in our network’s learned embedding space $\mathbb{R}^{D_{\gamma}}$ from value $\boldsymbol{\gamma}_{a}$ to zero and then to $\boldsymbol{\gamma}_{b}$. Bottom: A geometric interpolation in junction space $\mathbf{g}\in{\cal G}$ that passes through $\mathbf{g}_{0}=(\mathbf{0},0,\nicefrac{{1}}{{3}}\cdot\mathbf{1})$. The embedding has learned to be smooth and have an intuitive zero.
S4.3  Evolution of the Outputs
Figure S5 shows an example of how the distance map $\bar{d}[n]$ evolves during refinement. Specifically, we visualize the result of slicing $b_{k}(x;\mathbf{g})$, which we then apply a non-linearity to in order to amplify the less prominent boundaries. We see that early iterations are exploratory and unstructured, and that later iterations reach agreement.
Fig. S5: Evolution of boundaries during iterations, in reading order. Early iterations are exploratory and unstructured, while later iterations feature consistent per-patch boundaries, resulting in clean average boundary maps.
S4.4  Qualitative Results for Natural Images
@@
Fig. S6: Qualitative behavior of our model’s output boundaries $\bar{b}_{\eta}[n]$ on noiseless natural images, compared to those of end-to-end models EDTER [19] and Pidinet [23] that are trained to match human annotations; and compared to two bottom-up methods that, like our model, are not trained to match human annotations: Canny [3], and Field of Junctions [25] with patch size $11$.
In Figures S7 and S6, we show how the model behaves on noiseless natural images that contain texture and recognizable objects. In particular, Figure S6 emphasizes how the boundary maps produced by our model qualitatively differ from those of many classical bottom-up edge-detectors and also from those of learned, end-to-end models that have been trained to match human annotations. The figure compares our output to that from Canny [3], Field of Junctions [25] with a patch size of $11$, Pidinet [23], and EDTER [19]. The latter two methods are trained on human annotated data, whereas the former two methods, like our model, are not. (Note that inputs for all models besides EDTER [19] were $300\times 400$. Input to EDTER was down-sampled to $225\times 300$ due to its input size constraint.)
We find that our model produces finer structures than the end-to-end learned models [7, 19] because it is trained to only use local spatial averages of color as its cue for boundaries and grouping. It does not include mechanisms for grouping based on local texture statistics, nor based on non-local shape and appearance patterns that have semantic meaning to humans. Compared to the bottom-up methods of Canny [3] and Field of Junctions [25], our model has the advantage of automatically adapting the sizes of its output structures across the image plane, through its prediction of windowing field $\mathbf{p}[k]$. In contrast, the Field of Junctions and Canny both operate at a single pre-determined choice of local size, so they tend to oversegment some places while undersegmenting others.
@@
Fig. S7: Qualitative behavior of our model on noiseless natural images. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}[n]$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
S5 Additional Examples for Low-light Images
@@
Figure S8 shows examples of applying our model to indoor images taken by an iPhone XS in low light conditions.
@@
Figure S9 provides additional comparisons for a sample of varying-noise images from the ELD dataset [26]. When detecting boundaries at low signal-to-noise ratios, it is difficult to accurately discern finer structures as the noise level increases. Some algorithms, such as Field of Junctions [25], have tunable parameters such as patch-size that provide control over the level of detection. A small patchsize allows recovering fine structures in lower noise situations, but it causes many false positive boundaries at high noise levels. Conversely, a large patchsize provides more resilience to noise but has not ability to recover fine structure at all. Our model reduces the severity of this trade-off by automatically adapting its local windowing functions in ways that have learned to account for both the amount of noise and the local geometry of the underlying boundaries.
@@
In Figure S9 we see that our model is able to capture the double-contour shape of the curved, thin black bars, and that it continues to resolve them as the noise level increases, more than the other low-level methods. We also note that only the low-level models resolve this level of detail in the first place: The models trained on human annotations—UAED [28], EDTER, HED, Pidinet, and Structured Forests—miss the double contour entirely, estimating instead a single thick curve. We emphasize again that a user can adjust the behavior of Canny and Field of Junctions by tuning their local size parameters, either the filter size for Canny or the patchsize for Field of Junctions. Increasing the local size improves their resilience to noise but reduces their spatial precision. Neither system provides the ability to estimate fine grained details and withstand noise, like our model does.
@@
Figure S10 contains additional examples of images cropped from the ELD dataset. Here we include examples with even higher levels of noise to show the complete degradation of our algorithm and others.
Fig. S8: Visualization of our model’s output for low-light images captured by an iPhone XS. From left to right: Input image $\mathbf{f}[n]$, output distance map $\bar{d}[n]$, output boundary map $\bar{b}_{\eta}[n]$ with $\eta=0.7$, and output boundary-smoothed features $\bar{\mathbf{f}}[n]$.
Fig. S9: Qualitative comparison between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, for a crop from the ELD dataset under increasing amounts of photographic noise. We compare to end-to-end models that are trained to match human annotations (UAED [28], EDTER [19], HED [27], Pidinet [23], and Structured Edges (SE) [7]) in addition to low-level models that are not (Canny [3], gPb [14], and Field of Junctions (FoJ-17) [25]) with patch size 17.
Fig. S10:
Fig. S10: (cont.)
Fig. S10: (cont.)
@@
Fig. S10: (cont.) Additional qualitative comparisons between our model’s output boundaries $\bar{b}_{\eta}[n]$ and those of other methods, using crops from the ELD dataset under increasing amounts of photographic noise, including very high levels of noise.
S6 Additional Uses of Our Model
@@
S6.1  Color-based Depth Completion
Figure S11 shows an example of using our model for simple hole-filling in the depth channels of RGBD images from the Middlebury Stereo Datasets [20, 21]. We run our model on the RGB channels, and then for each pixel $n$ that has a missing depth value, we use our model’s output local attention kernels $a_{n}(x)$ to fill in that pixel’s value using an attention-weighted average of the observed depth values around it. This simple algorithm can be applied whenever the hole sizes are smaller than the maximum diameter of our attention maps, which is $34\times 34$ in our current implementation).
Fig. S11: Using our model for depth completion in RGBD images. Left: Input RGB channels. Middle: Input depth channel, with dark blue indicating missing values. Right: Completed depth using our model’s output attention kernels.
S6.2  Application: Photo Stylization
Figure S12 shows examples of using our model’s output for image stylization, by superimposing an inverted copy of the output boundary map $\bar{b}_{\eta}[n]$ onto the smoothed colors $\bar{\mathbf{f}}[n]$.
Fig. S12: Examples of stylized natural photographs, created by imposing our method’s output boundary map onto the output smoothed colors.